I am having trouble understanding functions. I am an on line student and my homework program had problems that sis not look quite like the one on a practice test i am taking. This is the question (I am sorry i am trying to follow the rules above but it is a 5 part question so only answer what you think would help me grasp this the most):
a graph accompanies this question with points in this order from right to left:
(-8,0)(-5,3)(0,-4)(5,0)(7,6)
Given the following function:
a) f(-5)=
b) f(0)=
c) What is the domain? {-8,7} (these I am somewhat comfortable with)
d) What is the range? {-4,6}
e) Find the value(s)of x such that f(x)=0

Found 2 solutions by stanbon, Edwin McCravy:
Answer by stanbon(75887)

NOTEl: I assume your points are in the order (x,y)
-------------------------
a graph accompanies this question with points in this order from right to left:
(-8,0)(-5,3)(0,-4)(5,0)(7,6)
Given the following function:
a) f(-5)= 3
------------
b) f(0)= -4
------------
c) What is the domain?:: {-8,-5,0,5,7}
---------------------------------------
d) What is the range? :: {0,3,-4,6}
----------------------------------------
e) Find the value(s)of x such that f(x)=0::x = -8 and x = 5
================================================================
Comment: Be careful. You may be wasting your time and money
taking an online course. Most people need to taught math.
===================================
Cheers,
Stan H.

Answer by Edwin McCravy(20059)

What you must do is remember that when you see this

(-8,0)(-5,3)(0,-4)(5,0)(7,6) ,

this is what it means:
 
#1.  (-8,0) means x=-8 and y=f(x)=f(-8)=0
#2.  (-5,3) means x=-5 and y=f(x)=f(-5)=3
#3.  (0,-4) means  x=0 and y=f(x)=f(0)=-4
#4.  (5,0) means  x=5 and y=f(x)=f(5)=0
#5.  (7,6) means  x=7 and y=f(x)=f(7)=6
 
a) f(-5)  

Looking above you can see from #2 that the answer is 3.

b) f(0)=

Looking above you can see from #3 that the answer is -4. 

c) What is the domain? 

Think of them as:

(-8,0)(-5,3)(0,-4)(5,0)(-7,6)

The domain is the set of all the red ones, the first ones, the x's:

domain = {-8,-5,0,5,7}

d) What is the range?

The range is the set of all the green ones, the second ones, the y's,
also called the f(x)'s:

{0,3,-4,0,6), but you shouldn't list 0 twice, so you
just have:

{0,3,-4,6)

But you also should put these in order from smallest to largest:

{0,-4,3,6)

e) Find the value(s)of x such that f(x)=0

Looking above at #1 and #4 you can see that there are two answers,
-8 and 5.

Edwin